OpenStudy (anonymous):

which of the following are the x and y intercepts of the equation 4x-2y=8? A.(3,0) and (0,4) B.(2,0) and (0,4) C.(2,0) and (0,-4)

OpenStudy (anonymous):

First put it in slope form -2y=-4x+8 y=2x-4

OpenStudy (anonymous):

To find the y intercept substitute 0 for x, so y=0-4 y=-4 So the y intercept would be (0,-4)

OpenStudy (anonymous):

To find the x intercept substitute 0 for y,so 0=2x-4 4=2x x=2 So the x intercept would be (2,0)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Your welcome :)